VIBRATIONAL PREDISSOCIATION SPECTROSCOPY OF THE (H_{2}O)_{n}^{−} n=6-21 CLUSTERS IN THE OH STRETCHING REGION: EVOLUTION OF THE EXCESS ELECTRON-BINDING SIGNATURE INTO THE INTERMEDIATE CLUSTER SIZE REGIME

\underline{JOSEPH C. BOPP}, JOSEPH R. ROSCIOLI, NATHAN I. HAMMER, MARK A. JOHNSON, Sterling Chemistry Laboratory, Yale University, PO Box 208107, New Haven, CT 06520.

We report vibrational predissociation spectra of the (H_{2}O)_{n}^{−} cluster ions in the OH stretching region to determine whether the spectral signature of the electron-binding motif identified in the smaller clusters continues to be important in the intermediate size regime (n=7-21). This signature consists of a redshifted doublet that dominates the OH stretching region, and has been traced primarily to the excitation of a single water molecule residing in a double H-bond acceptor (AA) binding site, oriented with both of its H atoms pointing toward the excess electron cloud. Strong absorption near the characteristic AA doublet is found to persist in the spectra of the larger clusters, but the pattern evolves into a broadened triplet around n=11. A single free OH feature associated with dangling hydrogen atoms on the cluster surface is observed to emerge for n=15, in sharp contrast to the multiplet pattern of unbonded OH stretches displayed by the H^{+}·(H_{2}O)_{n} clusters throughout the n=2-29 range.